# ICPO: Intrinsic Confidence-Driven Group Relative Preference Optimization for Efficient Reinforcement Learning

### v4

@@ 4.1 Baselines and Setup
We conduct experiments on Gemma2 Gemma2-2024, Llama3.1  llama3-2024 and Qwen2.5 Qwen2.5-2024 series models for fair comparison with existing methods. Unless otherwise specified, experiments are conducted on Qwen2.5-7B-Base. A comprehensive performance evaluation of ICPO method was carried out using three mathematical reasoning benchmarks, namely MATH-500 MATH-500-2024, Minerva Minerva-2022, and AIME24, as well as four general-domain benchmarks, including MMLU-Pro Mmlu-pro-2024, GPQA GPQA-2023, TheoremQA TheoremQA-2023, and WebInstruct. The baseline models include: (1) Base and Instruct models; (2) PRIME Ganqu-CuiLifan-Yuan-2025; (3) SimpleRL-Zoo SimpleRL-Zoo-2025; (4) Oat-Zero Zichen-Liu-2025; (5) TTRL TTRL-2025; (6) General Reasoner General-Reasoner-2025; (7) RLPR RLPR-2025; (8) INTUITOR Xuandong-Zhao-2025; and (9) CDE CDE-2025. Detailed descriptions of datasets and experimental setups are elaborated upon in the Appendix.
@@ 4.2 Main Results
The main results are presented in Table 1 while the training dynamic is presented in Figure 3. The key observations are as follows:
(1) Intrinsic confidence-driven mechanism has substantially enhanced reasoning performance in general domains. Compared to RLVR trained using the vanilla GRPO, we observed greater improvements in general reasoning performance across the Gemma, Llama, and Qwen models, with average increases of 1.4, 2.7, and 2.5 points, respectively. Furthermore, ICPO also demonstrated a significant advantage over other baseline methods.
(2) ICPO demonstrates exceptional performance in mathematical reasoning scenarios, surpassing both PRIME and SimpleRL-Zoo across three mathematical benchmarks. Notably, even without training on mathematical models, ICPO still demonstrates a significant enhancement in mathematical reasoning capability, comparable to frameworks specifically tailored for math reasoning.
Figure 3: Comparison of Avg accuracy on seven benchmarks and entropy over training of vanilla GRPO (Baseline) and ICPO (Our methods).
@@
Table 2: Performance on coarse-grained reward settings.
@@ 4.3 Analysis of ICPO
(1) Beyond Binary Rewards: Providing Continuous Optimization Signals. We filtered the training data, keeping only samples where all instances were entirely correct or entirely incorrect, and used them as coarse-grained reward training set. As shown in Table 2, compared to GRPO, ICPO still effectively explores under complete reward sparsity: it progressively identifies high-potential response paths and ultimately achieves stable performance gains. In contrast, GRPO, lacking fine-grained feedback, fails to pinpoint accurate policy update directions, resulting in stagnant or even deteriorating performance. These findings confirm that ICPO, through implicit contrastive modeling, effectively mitigates the exploration-exploitation dilemma in extreme coarse-grained reward settings, offering a viable optimization approach for scenarios without explicit preference labels.
@@ 4.4 Natural Scalability
ICPO exhibits scalability and can effectively adapt to models of varying scales. As shown in Table 3, with the growth of model parameters (from 3B to 7B, then to 14B), its performance keeps improving, fully reflecting ICPO’s inherent scalability. An interesting finding is that ICPO shows the greatest performance gain on 7B-scale model, with smaller improvements on 3B and 14B-scale models. We posit that this may stem from the non-linear relationship between model scale and confidence calibration: small models lack confidence, providing uncertain probability signals for many questions, while large models tend to be overconfident (ConfRAG-2025), assigning high confidence even to wrong answers. Consequently, both types of models exhibit relatively flat probability distributions, resulting in small in-group confidence differences.
@@
Notably, GRPO training on Qwen2.5-7B-Inst. shows sustained performance decline in later stages (as shown in Figure 5), whereas ICPO maintains stable exploration. We attribute this to the substantial performance improvement of the instruction-tuned model, which results in smaller relative advantages among different responses within groups. This further validating the robustness and anti-degradation benefits of intrinsic confidence-driven mechanism in coarse-grained reward scenarios.
@@ 4.5 Investigate the Effect of Intrinsic Bonus
Additionally, as previously indicated by research, entropy offers a crucial perspective for understanding exploration capabilities Ganqu-Cui-2025. A sharp decline or high-level fluctuation in entropy typically indicates insufficient model exploration or policy failure. Figure 2 (right) presents a comparative analysis of entropy dynamics between vanilla GRPO and ICPO. Firstly, the intrinsic confidence-driven mechanism effectively mitigates the phenomenon of entropy collapse, fully demonstrating its role in promoting exploration. Secondly, when comparing different decay schemes, the warmup decay approach yields a more stable entropy trajectory. This finding aligns with our understanding: prioritizing warmup and appropriately decaying reward weights ensures stable convergence while effectively supporting the exploration process.
@@ 5 Conclusion
Although ICPO can achieve effective learning in coarse-grained reward settings and mitigate the problem of noisy rewards, it still encounters limitations and challenges when rewards are completely inaccurate. The fundamental limitation is that ICPO inherently promotes in-depth learning of correct but low-probability samples while curbing overconfidence in incorrect yet high-probability ones. When rewards completely deviate from the true task objectives, this encouragement-suppression mechanism experiences a total reversal. Specifically, genuinely valid low-probability behaviors are incorrectly marked as erroneous samples by the inaccurate rewards, while truly ineffective or even detrimental behaviors are falsely portrayed as correct but low-probability high-quality samples and become the focal point of ICPO’s optimization, ultimately causing policy failure. We are currently actively exploring solutions to these issues and will present them in our future research.
ICPO training prompt
<|im_start|>system
A conversation between User and Assistant. The user asks a question, and the Assistant solves it. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think> <answer> answer here </answer>.
<|im_end|>
<|im_start|>user
{{question}}<|im_end|>
<|im_start|>assistant
Table 5: We adopt the training prompt of R1 DeepSeek-R1-2025 for ICPO.
@@ A.1 Training data
Our training data is identical to that employed in RLPR RLPR-2025, utilizing the high-quality reasoning question bank released by Ma et al. General-Reasoner-2025, which encompasses high-caliber reasoning problems across multiple complex domains. This training dataset excludes mathematics-related prompts to focus on general-domain reasoning and adopts a multi-stage filtering strategy to ensure its difficulty level: initially, history-related questions and those at primary/junior high school levels are removed to avoid overly simplistic or common-sense content; subsequently, only high-difficulty samples are retained based on reasoning scores assigned by GPT-4.1-mini. Ultimately, this process yields 77,687 focused and high-quality complex non-mathematical reasoning data entries.
@@ A.2 Evaluation data
  • MATH-500 MATH-500-2024 is a challenging benchmark specifically designed for evaluating the mathematical reasoning capabilities of LLMs. It comprises 500 high-quality, challenging mathematical problems, aiming to comprehensively assess a model’s ability to solve complex mathematical issues.
  • Minerva Minerva-2022, developed by Google DeepMind, is a high-quality dataset for training mathematical reasoning models. It aims to narrow the capability gap of LLMs in quantitative reasoning, empowering models with mathematical thinking and symbolic reasoning skills from basic calculations to advanced scientific research levels.
@@
  • MMLU-Pro Mmlu-pro-2024 is a widely-adopted multi-task language understanding benchmark, encompassing challenging reasoning-intensive questions across multiple domains. To strike a balance between evaluation efficiency and data diversity, we randomly sampled 1,000 prompts from this benchmark for our experiments.
  • GPQA GPQA-2023 encompasses graduate-level questions across multiple disciplines such as physics and chemistry. We employ the highest-quality GPQA-diamond subset for our evaluation.
  • TheoremQA TheoremQA-2023 is designed to evaluate a model’s ability to solve complex scientific problems by applying theorems. This benchmark comprises 800 high-quality questions, covering 350 theorems across fields such as mathematics and physics. We remove the 53 multimodal instructions.
  • WebInstruct: We derived a validation subset from WebInstruct General-Reasoner-2025 to serve as an easily evaluable benchmark tailored for medium-sized models. Despite its relatively lower difficulty, the dataset can still evaluate multidisciplinary reasoning skills. Adhering to the filtering settings introduced in RLPR RLPR-2025, we conducted our evaluation using 638 unique questions.
@@ B.1 Training Setup
Unless otherwise specified, all our experiments were conducted on the Qwen2.5-7B model. Following the majority of RLVR practices, we omitted the supervised fine-tuning process and directly performed post-training on the base model. In the main experiment, we controlled the output structure to extract parsable reasoning chains and answers by adjusting the prompt templates during both the training and validation phases. The specific prompt templates are detailed in Table 5.
In the experiments targeting the Gemma and Llama models, we adjusted the temperature parameters for both training and evaluation to 0.6. Following RLPR (RLPR-2025) guidelines, we removed <think> segments from templates to prevent a decline in generation quality. We observed that rule-based scoring scripts introduced errors in benchmark tests containing non-multiple-choice formatted questions. To address this issue, we deployed a Qwen2.5-7B-Instruct model server for evaluation and additionally utilized Qwen2.5-72B-Instruct to handle more complex benchmarks, such as TheoremQA and Minerva.
In the experiment focusing on coarse-grained reward scenarios, we employed Qwen2.5-7B model to conduct a single-epoch training session on the training set. We retained the original data where all responses within each group were either uniformly correct or uniformly incorrect, while filtering out those data instances where responses within the same group exhibited discrepancies in rewards. Subsequently, we utilized the retained data as coarse-grained reward training set to further train Qwen2.5-7B model, thereby validating the efficacy of ICPO in coarse-grained reward settings.
@@ B.2 Parameters Setup
Appendix C Theoretical Explanation

### v5

@@ 4.1 Baselines and Setup
We conduct experiments on Gemma2 Riviere et al. (2024), Llama3.1  Grattafiori et al. (2024) and Qwen2.5 Yang et al. (2024) series models for fair comparison with existing methods. Unless otherwise specified, experiments are conducted on Qwen2.5-7B-Base. A comprehensive performance evaluation of ICPO method was carried out using three mathematical reasoning benchmarks, namely MATH-500 Cobbe et al. (2021), Minerva Lewkowycz et al. (2022), and AIME24, as well as four general-domain benchmarks, including MMLU-Pro Wang et al. (2024b), GPQA Rein et al. (2023), TheoremQA Chen et al. (2023), and WebInstruct. The baseline models include: (1) Base and Instruct models; (2) PRIME Cui et al. (2025a); (3) SimpleRL-Zoo Zeng et al. (2025); (4) Oat-Zero Liu et al. (2025); (5) TTRL Zuo et al. (2025); (6) General Reasoner Ma et al. (2025); (7) RLPR Yu et al. (2025b); (8) INTUITOR Zhao et al. (2025); and (9) CDE Dai et al. (2025). Detailed descriptions of datasets and experimental setups are elaborated upon in Appendix A and B.
@@ 4.2 Main Results
Main results and training dynamic are presented in Table 1 and Figure 3. Key findings are as follows:
(1) Intrinsic confidence-driven mechanism has substantially enhanced reasoning performance in general domains. Compared to RLVR trained using the vanilla GRPO, we observed greater improvements in general reasoning performance across the Gemma, Llama, and Qwen models, with average increases of 1.5, 2.8, and 2.5 points, respectively. Furthermore, ICPO also demonstrated a significant advantage over other baseline methods.
(2) ICPO demonstrates exceptional performance in mathematical reasoning scenarios, surpassing both PRIME and SimpleRL-Zoo across three mathematical benchmarks. Notably, even without training on mathematical models, ICPO still demonstrates a significant enhancement in mathematical reasoning capability, comparable to methods specifically trained on mathematical models.
Figure 3: Comparison of Avg accuracy on seven benchmarks and entropy of GRPO (Baseline) and ICPO (Our).
@@
Table 2: Performance on sparse reward settings.
@@ 4.3 Analysis of ICPO
(1) Beyond Binary Rewards: Providing Continuous Optimization Signals. We filtered the training data, keeping only samples where all instances were entirely correct or entirely incorrect, and used them as extremely sparse reward training set. As shown in Table 2, compared to GRPO, ICPO still effectively explores under complete reward sparsity: it progressively identifies high-potential response paths and ultimately achieves stable performance gains. In contrast, GRPO, lacking fine-grained feedback, fails to pinpoint accurate policy update directions, resulting in stagnant or even deteriorating performance. These findings confirm that ICPO, through implicit contrastive modeling, effectively mitigates the exploration-exploitation dilemma in extremely sparse reward settings, offering a viable optimization approach for scenarios without explicit preference labels.
@@ 4.4 Natural Scalability
ICPO exhibits scalability and can effectively adapt to models of varying scales. As shown in Table 3, with the growth of model parameters (from 3B to 7B, then to 14B), its performance keeps improving, fully reflecting ICPO’s inherent scalability. An interesting finding is that ICPO shows the greatest performance gain on 7B-scale model, with smaller improvements on 3B and 14B-scale models. We posit that this may stem from the non-linear relationship between model scale and confidence calibration: small models lack confidence, providing uncertain probability signals for many questions, while large models tend to be overconfident (Huang et al., 2025), assigning high confidence even to wrong answers. Consequently, both types of models exhibit relatively flat probability distributions, resulting in small in-group confidence differences.
@@
Notably, GRPO training on Qwen2.5-7B-Inst. shows sustained performance decline in later stages (as shown in Figure 5), whereas ICPO maintains stable exploration. We attribute this to the substantial performance improvement of the instruction-tuned model, which results in smaller or even zero relative advantages among responses within groups. This further validating the robustness and anti-degradation benefits of intrinsic confidence-driven mechanism in extremely sparse reward scenarios.
@@ 4.5 Investigate the Effect of Intrinsic Bonus
Additionally, as previously indicated by research, entropy offers a crucial perspective for understanding exploration capabilities Cui et al. (2025b). A sharp decline or high-level fluctuation in entropy typically indicates insufficient model exploration or policy failure. Figure 2 (right) presents a comparative analysis of entropy dynamics between vanilla GRPO and ICPO. Firstly, the intrinsic confidence-driven mechanism effectively mitigates the phenomenon of entropy collapse, fully demonstrating its role in promoting exploration. Secondly, when comparing different decay schemes, the warmup decay approach yields a more stable entropy trajectory. This finding aligns with our understanding: prioritizing warmup and appropriately decaying reward weights ensures stable convergence while effectively supporting the exploration process.
@@ 5 Conclusion
Although ICPO can achieve effective learning in sparse reward settings and mitigate the problem of noisy rewards, it still encounters limitations and challenges when rewards are completely inaccurate. The fundamental limitation is that ICPO inherently promotes in-depth learning of correct but low-probability samples while curbing overconfidence in incorrect yet high-probability ones. When rewards completely deviate from the true task objectives, this encouragement-suppression mechanism experiences a total reversal. Specifically, genuinely valid low-probability behaviors are incorrectly marked as erroneous samples by the inaccurate rewards, while truly ineffective or even detrimental behaviors are falsely portrayed as correct but low-probability high-quality samples and become the focal point of ICPO’s optimization, ultimately causing policy failure. We are currently actively exploring solutions to these issues and will present them in our future research.
@@ A.1 Training data
Our training data is identical to that employed in RLPR Yu et al. (2025b), utilizing the high-quality reasoning question bank released by Ma et al. Ma et al. (2025), which encompasses high-caliber reasoning problems across multiple complex domains. This training dataset excludes mathematics-related prompts to focus on general-domain reasoning and adopts a multi-stage filtering strategy to ensure its difficulty level: initially, history-related questions and those at primary/junior high school levels are removed to avoid overly simplistic or common-sense content; subsequently, only high-difficulty samples are retained based on reasoning scores assigned by GPT-4.1-mini. Ultimately, this process yields 77,687 focused and high-quality complex non-mathematical reasoning data entries.
@@ A.2 Evaluation data
  • MATH-500 Cobbe et al. (2021) is a challenging benchmark specifically designed for evaluating the mathematical reasoning capabilities of LLMs. It comprises 500 high-quality, challenging mathematical problems, aiming to comprehensively assess a model’s ability to solve complex mathematical issues.
  • Minerva Lewkowycz et al. (2022), developed by Google DeepMind, is a high-quality dataset for training mathematical reasoning models. It aims to narrow the capability gap of LLMs in quantitative reasoning, empowering models with mathematical thinking and symbolic reasoning skills from basic calculations to advanced scientific research levels.
@@
  • MMLU-Pro Wang et al. (2024b) is a widely-adopted multi-task language understanding benchmark, encompassing challenging reasoning-intensive questions across multiple domains. To strike a balance between evaluation efficiency and data diversity, we randomly sampled 1,000 prompts from this benchmark for our experiments.
  • GPQA Rein et al. (2023) encompasses graduate-level questions across multiple disciplines such as physics and chemistry. We employ the highest-quality GPQA-diamond subset for our evaluation.
  • TheoremQA Chen et al. (2023) is designed to evaluate a model’s ability to solve complex scientific problems by applying theorems. This benchmark comprises 800 high-quality questions, covering 350 theorems across fields such as mathematics and physics. We remove the 53 multimodal instructions.
  • WebInstruct: We derived a validation subset from WebInstruct Ma et al. (2025) to serve as an easily evaluable benchmark tailored for medium-sized models. Despite its relatively lower difficulty, the dataset can still evaluate multidisciplinary reasoning skills. Adhering to the filtering settings introduced in RLPR Yu et al. (2025b), we conducted our evaluation using 638 unique questions.
@@ B.1 Training Setup
Unless otherwise specified, all our experiments were conducted on the Qwen2.5-7B model. Following the majority of RLVR practices, we omitted the supervised fine-tuning process and directly performed post-training on the base model. In the main experiment, we controlled the output structure to extract parsable reasoning chains and answers by adjusting the prompt templates during both the training and validation phases. The specific prompt templates are detailed in Table 5. All experimental results presented in this paper are averaged over multiple trials.
In the experiments targeting the Gemma and Llama models, we adjusted the temperature parameters for both training and evaluation to 0.6. Following RLPR (Yu et al., 2025b) guidelines, we removed <think> segments from templates to prevent a decline in generation quality. We observed that rule-based scoring scripts introduced errors in benchmark tests containing non-multiple-choice formatted questions. To address this issue, we deployed a Qwen2.5-7B-Instruct model server for evaluation and additionally utilized Qwen2.5-72B-Instruct to handle more complex benchmarks, such as TheoremQA and Minerva.
In the experiment focusing on sparse reward scenarios, we employed Qwen2.5-7B model to conduct a single-epoch training session on the training set. We retained the original data where all responses within each group were either uniformly correct or uniformly incorrect, while filtering out those data instances where responses within the same group exhibited discrepancies in rewards. Subsequently, we utilized the retained data as extremely sparse reward training set to further train Qwen2.5-7B model, thereby validating the efficacy of ICPO in extremely sparse reward settings.
@@
Figure 6: Log-probability distribution of output responses from models of different scales. From left to right are the Qwen2.5-Base series, Qwen2.5-Instruct series, and Qwen3-Base series.
@@ B.2 Parameters Setup
Appendix C Hyperparameter Search Experiment
We conducted additional hyperparameter search experiments focusing on the following hyperparameters: (1) Selection method of parameter $\delta$; (2) Impact of $\tau$ on model performance; and (3) Analysis of the weight adjustment approach for $\omega$.
C.1 Selection method of parameter $\delta$
Table 7 presents a performance comparison of $\delta$ across models of different types and scales. Through analysis of the experimental results, we observe that the optimal value of $\delta$ varies across models of different scales. However, for models of different types (Qwen2.5-Base, Qwen2.5-Instruct, and Qwen3-Base), the optimal values of $\delta$ are approximately the same under the same scale.
Model GRPO ICPO
$\delta=0.2$ $\delta=0.3$ $\delta=0.4$ $\delta=0.5$ $\delta=0.6$
Qwen2.5-Base Models
Qwen2.5-3B-Base 40.10 40.30 40.88 40.70 41.10 40.85
Qwen2.5-7B-Base 52.20 52.03 54.23 54.90 53.65 52.90
Qwen2.5-14B-Base 57.05 57.63 57.85 57.24 57.30 56.45
Qwen2.5-Instruct Models
Qwen2.5-3B-Instruct 42.40 42.15 43.33 43.67 44.10 43.90
Qwen2.5-7B-Instruct 51.55 53.65 55.47 55.60 54.70 51.90
Qwen2.5-14B-Instruct 60.01 61.05 61.40 60.38 59.70 60.13
Qwen3-Base Models
Qwen3-4B-Base 49.55 52.20 53.35 54.05 53.90 52.67
Qwen3-8B-Base 57.30 59.27 59.15 59.93 58.80 57.33
Qwen3-14B-Base 58.93 59.70 60.57 59.48 59.20 58.75
Table 7: The impact of parameter $\delta$ on performance across models of different scales (3B, 7B and 14B), types (Base and Instruct) and versions (Qwen2.5 and Qwen3), where the performance is the average across seven benchmarks.
We analyze that the possible reason lies in the fact that the probability distributions of model outputs under the same scale are roughly similar, even when the model types differ. To verify this hypothesis, we conducted a statistical analysis of the logarithmic probability distributions output by different models and plotted scatter diagrams, as illustrated in Figure 6. The experimental results robustly validate our hypothesis and reveal a key insight: Smaller models exhibit low confidence, providing uncertain probabilistic signals for many queries, whereas larger models tend toward overconfidence, assigning high probabilities even to incorrect answers. Models of moderate scale (4B-8B parameters) demonstrate optimal consistency in response probability distributions. Moreover, for models of comparable scales, their output probability distributions exhibit substantial similarity. Consequently, the configuration of parameter $\delta$ demonstrates general applicability across different model types. Specifically, we recommend setting $\delta=0.5$ for models with fewer than 4B parameters, $\delta=0.4$ for models ranging from 4B to 8B parameters, and $\delta=0.3$ for models exceeding 8B parameters.
C.2 Impact of parameter $\tau$ on performance
The core function of parameter $\tau$ is to regulate the weight allocation of intrinsic confidence rewards within the overall reward, thereby determining its proportion of influence on the policy update process. If $\tau$ is set too large, the policy update will be excessively dominated by intrinsic preferences, making it difficult to ensure the performance effectiveness of the policy in real-world task scenarios. Conversely, if $\tau$ is set too small, the guiding role of intrinsic preferences in the optimization process will be weakened, failing to effectively calibrate the direction of policy updates. We set the values of parameter $\tau$ to 1.0, 2.0, 3.0, 4.0, and 5.0 respectively, and observed the changes in the performance of ICPO under different settings. The experimental results are shown in Figure 7.
Figure 7: Impact of parameter $\tau$ on performance.
From the results, we observe that the model achieves optimal performance when the parameter $\tau=2.0$, with $\tau=3.0$ yielding the second-best yet still competitive performance. In contrast, setting $\tau=1.0$ leads to a significant degradation in both policy entropy and overall performance, indicating that over-reliance on intrinsic preferences for guiding policy updates destabilizes training. When $\tau$ is further increased to 4.0 or 5.0, the model exhibits rapid performance gains in the early stages of training but subsequently plateaus, accompanied by a sharp decline in policy entropy. This suggests that excessively downweighting the influence of intrinsic preferences during policy updates undermines the agent’s exploratory capacity. Collectively, these findings demonstrate that appropriately leveraging intrinsic preferences to guide policy updates helps mitigate entropy collapse and effectively preserves policy exploration.
C.3 Analysis of Weight Adjustment Strategies
Table 8 presents the hyperparameter search experiments for the weight adjustment strategies across different models. Figure 8 illustrates the parameter search experiment for the warm-up turning point. From these, we observe that: (1) Weight decay mechanism is of critical importance. For Qwen2.5-7B-Base, Qwen2.5-7B-Instruct, and Qwen3-4B-Base, all decay schemes outperform the no-decay baseline, enabling a smooth transition from exploration to exploitation. (2) For the Qwen2.5-7B-Base and Qwen3-4B-Base models, warm-up exploration in the early stages is crucial. However, for the Qwen2.5-7B-Instruct model, the impact of warm-up exploration on performance improvement is not significant. (3) For the Qwen2.5-7B-Base model, the warm-up decay weight adjustment method achieves optimal performance when the warm-up turning point is set to 0.4. Both arriving at the warm-up turning time point earlier or later will exert a certain impact on the exploration-exploitation effectiveness of the strategy.
Figure 8: Parameter search for warm-up turning point.
Model MMLUPro GPQA TheoremQA WebInst. MATH500 Minerva AIME24 All
Avg@2 Avg@4 Avg@2 Avg@2 Avg@2 Avg@2 Avg@16 -
Bonus Weight Adjustment Schedules
Qwen2.5-7B-Base-GRPO 55.1 36.2 52.2 75.3 76.5 54.9 17.7 52.20
$\Rightarrow$ $\omega$ No decay 54.8 35.8 54.0 77.0 76.0 55.9 16.7 52.89
$\Rightarrow$ $\omega$ Linear decay 56.0 39.0 53.5 75.8 76.2 55.4 20.0 53.70
$\Rightarrow$ $\omega$ Warmup retention 57.0 38.7 53.5 76.2 76.5 55.4 20.0 53.90
$\Rightarrow$ $\omega$ Warmup decay 57.6 39.4 54.0 77.8 76.2 56.2 23.3 54.90
Qwen2.5-7B-Inst.-GRPO 52.0 40.1 56.3 63.2 77.0 55.8 17.7 51.73
$\Rightarrow$ $\omega$ No decay 57.7 42.2 52.1 76.8 74.5 52.6 20.0 53.70
$\Rightarrow$ $\omega$ Linear decay 58.3 43.8 54.7 77.3 77.0 54.2 23.3 55.51
$\Rightarrow$ $\omega$ Warmup retention 57.4 42.0 54.3 75.2 75.7 53.5 20.0 54.01
$\Rightarrow$ $\omega$ Warmup decay 58.1 42.6 55.4 74.9 77.8 55.8 26.7 55.90
Qwen3-4B-Base-GRPO 57.3 41.5 53.3 65.8 78.5 54.7 16.7 52.54
$\Rightarrow$ $\omega$ No decay 53.6 40.8 54.5 65.6 79.2 54.8 16.7 52.17
$\Rightarrow$ $\omega$ Linear decay 58.1 43.2 56.1 71.4 82.2 55.7 20.0 55.24
$\Rightarrow$ $\omega$ Warmup retention 57.5 43.5 55.7 68.2 80.8 56.9 20.0 54.66
$\Rightarrow$ $\omega$ Warmup decay 59.5 43.9 56.8 73.1 81.7 57.8 20.0 56.11
Table 8: Performance comparison of ICPO under different weight adjustment schemes for intrinsic confidence reward across different models. The weight adjustment schemes adhere to those illustrated in Figure 2.
(a) 0% of entire training process
(b) 20% of entire training process
(c) 40% of entire training process
(d) 60% of entire training process
(e) 80% of entire training process
(f) 100% of entire training process
Figure 9: The distribution of generation probabilities for correct responses and incorrect responses across different training phases throughout the entire training process on the Qwen2.5-7B-Base model.
These experimental results are precisely consistent with the key insights we proposed in Section 3.4: (1) In the early stages of training, a high-entropy environment makes it difficult to determine whether low-probability responses are worth learning. As training progresses, the entropy gradually decreases, and correct yet novel low-probability responses begin to emerge, thereby gradually enhancing the motivation to learn these responses. (2) In the later stages of training, when most novel samples have already been mastered, the remaining low-probability responses may merely constitute noise or valueless signals; thus, reducing the influence of intrinsic signals helps stabilize policy learning. (3) The period during which the warm-up phase sharply rises and reaches its peak turning point should align with the period when the entropy gradually decreases to an appropriate level while still retaining active exploration behavior.
For the Qwen2.5-7B-Instruct model, there is little difference in performance between warm-up exploration and directly applying weight decay. The primary reason lies in the fact that this model, having undergone instruction fine-tuning training on top of a pre-trained model, starts with a relatively low entropy level in the initial training phase (absent of a high-entropy environment), thus allowing for selective warm-up exploration. In contrast, for models like Qwen2.5-7B-Base and Qwen3-4B-Base, which are solely pre-trained, the presence of a high-entropy environment in the early training stages makes it more effective to gradually enhance their learning motivation for correct yet low-probability responses.
(a) 0.1 random reward noise
(b) 0.2 random reward noise
(c) 0.4 random reward noise
(d) 0.5 random reward noise
(e) 0.6 random reward noise
(f) 0.7 random reward noise
(g) 0.8 random reward noise
(h) 0.9 random reward noise
Figure 10: Comparison of ICPO and GRPO Performance under Different Intensities of Random Noise Injection during the Training Process Based on the Qwen2.5-7B-Base Model.
(a) 0.1 random reward noise
(b) 0.2 random reward noise
(c) 0.4 random reward noise
(d) 0.5 random reward noise
(e) 0.6 random reward noise
(f) 0.7 random reward noise
(g) 0.8 random reward noise
(h) 0.9 random reward noise
Figure 11: Comparison of ICPO and GRPO Entropy under Different Intensities of Random Noise Injection during the Training Process Based on the Qwen2.5-7B-Base Model.
To validate our key insights, we plotted the probability distributions of generating correct and incorrect responses across different training periods throughout the entire training process, as shown in Figure 9. It can be observed that: (1) In the very initial stage of training, there are extremely few correct yet low-probability responses, with the majority of low-probability responses being incorrect ones, making it impossible to effectively distinguish novel responses with learning value. (2) As training progresses to the 20%-40% stage, correct yet low-probability responses increase significantly, revealing a large number of novel samples with high learning value. (3) In the later stages of training, the number of correct yet low-probability samples returns to a relatively low level, with most low-probability samples now being incorrect ones (data noise or overly complex problems), whose learning value diminishes. This is highly consistent with our proposed viewpoints and also provides an intuitive explanation for why the warm-up decay weight adjustment scheme is effective.
Appendix D Ablation Study on Noisy Rewards
To further investigate the performance of ICPO in scenarios with noisy rewards, we varied the extent of random noise injected into the rewards. In the noise reward experiment in Section 4.3, we randomly selected 40% of the data during each update iteration and randomly added or subtracted 0.3 from the rewards of this portion of data as noise. Here, we further explored the comparative performance of ICPO and GRPO when the random noise levels were set to 0.1, 0.2, 0.4, 0.5, 0.6, 0.7, 0.8, and 0.9. The training dynamics are illustrated in Figures 10 and 11. It can be observed that when the random noise ranges from 0.1 to 0.5, ICPO can still effectively learn through its intrinsic confidence-driven mechanism, maintaining positive reinforcement for correct yet low-probability responses and effectively suppressing incorrect yet high-probability responses when external rewards are distorted, namely the two scenarios mentioned in Appendix E.3.
However, when the random noise in rewards exceeds 0.6, ICPO’s noise resistance reaches a turning point, which aligns precisely with the viewpoint we mentioned in the section on limitations: when rewards completely deviate from the true objectives of the task, this "encouragement-suppression" mechanism undergoes a complete reverse shift. Originally genuine and effective low-probability behaviors are mistakenly labeled as incorrect samples by the erroneous rewards, while those genuinely ineffective or even harmful behaviors are mispackaged by the erroneous rewards as correct yet low-probability high-quality samples, becoming the primary focus of ICPO’s optimization efforts and ultimately leading to policy collapse. This also represents a major direction for future optimization of ICPO.
Appendix E Theoretical Explanation